### v1

# Sample and Map from a Single Convex Potential: Generation using Conjugate Moment Measures

Abstract
A common approach to generative modeling is to split model-fitting into two blocks: define first how to sample noise (e.g. Gaussian) and choose next what to do with it (e.g. using a single map or flows). We explore in this work an alternative route that ties sampling and mapping. We find inspiration in moment measures (Cordero-Erausquin and Klartag, 2015), a result that states that for any measure $\rho$ supported on a compact convex set of $\mathbb{R}^{d}$, there exists a unique convex potential $u$ such that $\rho=\nabla u\,\sharp\,e^{-u}$. While this does seem to tie effectively sampling (from log-concave distribution $e^{-u}$) and action (pushing particles through $\nabla u$), we observe on simple examples (e.g., Gaussians or 1D distributions) that this choice is ill-suited for practical tasks. We study an alternative factorization, where $\rho$ is factorized as $\nabla w^{*}\,\sharp\,e^{-w}$, where $w^{*}$ is the convex conjugate of $w$. We call this approach conjugate moment measures, and show far more intuitive results on these examples. Because $\nabla w^{*}$ is the Monge map between the log-concave distribution $e^{-w}$ and $\rho$, we rely on optimal transport solvers to propose an algorithm to recover $w$ from samples of $\rho$, and parameterize $w$ as an input-convex neural network.
1 Introduction
A decade after the introduction of GAN  (Goodfellow et al., 2014) and VAE (Kingma and Welling, 2014), the field of generative modeling has grown into one of the most important areas of research in machine learning. Both canonical approaches follow the template of learning a transformation that can map random codes to meaningful data. These transformations can be learned in supervised manner, as in the dimensionality reduction pipeline advocated in VAEs, or distributionally as advocated in the purely generative literature. In the latter, the variety of such transforms has gained remarkably in both complexity, using increasingly creative inductive biases, from Neural-ODEs(Chen et al., 2018; Grathwohl et al., 2018), diffusions (Song et al., 2020; Ho et al., 2020), optimal transport (Korotin et al., 2020) to flow-matching (Lipman et al., 2023; Tong et al., 2023; Pooladian et al., 2023).
Sample first, move next. All of these approaches share, however, a standard Gaussian multivariate distribution to generate noise/codes. For both GANs and VAEs, that choice is usually made because of its simplicity, but also, in the case of diffusion models, because Gaussian distributions can be recovered quickly with suitable stochastic processes (Ohrstein-Uhlenbeck). Other works have considered optimizing prior noise distributions (Tomczak and Welling, 2018; Lee et al., 2021; Liang et al., 2022) but still do so in a two step approach where mappings are estimated independently. We propose in this work a new approach where sampling and transforms are not treated separately, but seen as two facets of the same single object.
Tying sampling and action. Our work looks to moment measures Cordero-Erausquin and Klartag (2015) to find inspiration for a factorization that ties both sampling and action.  (Cordero-Erausquin and Klartag, 2015) proved that for any sufficiently regular probability measure $\rho$, one can find an essentially-continuous, convex potential $u$ such that $\rho=\nabla u\,\sharp\,e^{-u}$: essentially, any probability distribution can be recovered by first sampling points from a log-concave distribution $e^{-u}$, and moving them with $\nabla u$ where, remarkably, the same single convex potential $u$ is used twice.
Our Contributions. We build on the work of  Cordero-Erausquin and Klartag (2015) to propose an alternative path to build generative models, that we call conjugate moment measures.
  • After recalling in detail the contribution of  Cordero-Erausquin and Klartag (2015) in §2.1, we argue in §3.1 that the moment measure factorization may not be suitable for practical tasks. For instance, in the case of Gaussian distributions, we show that if $\rho$ is Gaussian with variance $\Sigma$, the corresponding log-concave distribution $e^{-u}$ is Gaussian with variance $\Sigma^{-1}$, amplifying beyond necessary any minor degeneracy in $\Sigma$. Similarly, for univariate $\rho$, we found that the spread of $e^{-u}$ is inversely proportional to that of $\rho$. While these results are not surprising from a theoretical perspective, they suggest a different strategy to estimate a tied sample/action factorization.
  • We propose instead in §3.2 a new factorization that is conjugate to that of Cordero-Erausquin and Klartag (2015): We show that any absolutely continuous probability distribution $\rho$ supported on a compact convex set can be written as $\rho=\nabla w^{*}\,\sharp\,e^{-w}$, where $w^{*}$ is the convex conjugate of $w$. Importantly, in our factorization the convex potential $w$ is used to map $\rho$ to $e^{-w}$, since one has equivalently that $\nabla w\,\sharp\,\rho=e^{-w}$. Note that evaluating $w^{*}$, the convex conjugate of $w$, only requires solving a convex optimization problem, since $w$ is assumed to be convex.
  • We explore how to infer the potential $w$ arising from the conjugate factorization above using only samples from $\rho$, by relying on optimal transport (OT) theory (Santambrogio, 2015). Indeed, since the condition $\rho=\nabla w^{*}\,\sharp\,e^{-w}$ is equivalent to stating that $w^{*}$ is the Brenier potential linking $e^{-w}$ to $\rho$, we can parameterize $w$ as an input-convex neural network $w_{\theta}$ whose gradient is an OT map. We provide illustrations on simple generative examples.
2 Background
We present neural approaches to solve the OT problem and introduce moment measures.
2.1 Optimal Transport
Let $\mathcal{P}(\mathbb{R}^{d})$ be the set of Borel probability measures on $\mathbb{R}^{d}$. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$ and set $\Pi(\mu,\nu)$ to be the set of couplings between $\mu$ and $\nu$. We consider the primal OT problem with squared-Euclidean cost,
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\tfrac{1}{2}\|x-y\|^{2}\ \mathrm{d}\pi(x,y)\,.$ (1)
This problem admits a dual formulation:
$\begin{split}&\sup\limits_{(f,g)\in\textrm{L}^{1}(\mu)\times\textrm{L}^{1}(\nu)}\int_{\mathbb{R}^{d}}f(x)\mathrm{d}\mu(x)+\int_{\mathbb{R}^{d}}g(y)\mathrm{d}\nu(y),\\ &\textrm{s.t.}\;f(x)+g(y)\leq\tfrac{1}{2}\|x-y\|^{2},\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\,.\end{split}$ (2)
The functions $f$ and $g$ that maximize the expression in equation 2 are referred to as the Kantorovich potentials. Equation 1 can itself be seen as a relaxation of the Monge formulation of the OT problem,
$\inf_{\begin{subarray}{c}T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},\\ T\,\sharp\,\mu=\nu\end{subarray}}\int_{\mathbb{R}^{d}}\tfrac{1}{2}\|x-T(x)\|^{2}\mu(\mathrm{d}x)$ (3)
[…]
$\min_{\begin{subarray}{c}u\in\mathcal{C}(\mathbb{R}^{d})\\ u(0)=0\end{subarray}}\!\!\mathcal{T}_{u}(\mu,\nu):=\!\int_{\mathbb{R}^{d}}\!u(x)\,\mathrm{d}\mu(x)+\!\int_{\mathbb{R}^{d}}\!u^{*}(x)\,\mathrm{d}\nu(x)$ (4)
where $u^{*}(y):=\max_{x\in\mathbb{R}^{d}}\langle x,y\rangle-u(x)$ is the convex-conjugate of $u$ and noting that we lift any ambiguity on $u$ by selecting the unique potential such that $u(0)=0$. By denoting $\mathcal{T}(\mu,\nu)$ the optimal value and $\mathfrak{B}(\mu,\nu)$ the Brenier potential solving equation 4 above, this results in $\nu=\nabla\mathfrak{B}(\mu,\nu)\,\sharp\,\mu$.
Neural OT solvers. The goal of neural OT solvers is to estimate $\mathfrak{B}(\mu,\nu)$ using samples drawn from the source $\mu$ and the target distribution $\nu$. Makkuva et al. (2020); Korotin et al. (2020) proposed methods that leverage input convex neural networks (ICNN), originally introduced by Amos et al. (2017), to parameterize the potential $u$ as an ICNN. A key challenge in these approaches is managing the Legendre transform $u^{*}$. To overcome this, a surrogate network can be employed to approximate $u^{*}$ with recent advances by Amos (2023) improving these implementations through amortized optimization.
2.2 Moment Measures
For a given convex function $u$, the moment measure of $u$ is defined as the pushforward measure $\nabla u\,\sharp\,\mathfrak{P}_{u}$ where $\mathfrak{P}_{u}$ is the (Gibbs) log-concave probability measure with density
$\mathfrak{P}_{u}(x):=\frac{e^{-u(x)}}{\int_{\mathbb{R}^{d}}e^{-u(z)}\mathrm{d}z}\,.$ (5)
The moment measure of $u$ is in particular well defined when the quantity $\int_{\mathbb{R}^{d}}e^{-u(z)}\mathrm{d}z$ is positive and finite.
While any measure $\rho$ supported on a compact set $K$ is guaranteed to be the moment measure of some convex potential $u$ (Santambrogio, 2016), Cordero-Erausquin and Klartag showed that $u$ is often discontinuous at the boundary $\partial K$. In Cordero-Erausquin and Klartag (2015), they studied the variational problem
$\min_{u\in\mathcal{C}(\mathbb{R}^{d})}\int u^{*}\mathrm{d}\rho-\ln\left(\int e^{-u}\right)\,,$ (6)
and proved that if $\rho$ is a probability measure with a finite first moment, whose barycenter is at 00 and is not supported on a hyperplane, then it is the moment measure of an essentially continuous potential. Subsequently, Santambrogio (2016) derived the same result using a variational formulation over probability measures, given by
$\min_{\xi\in\mathcal{P}(\mathbb{R}^{d})}\{\mathcal{E}(\xi)+\mathcal{T}(\xi,\rho):\xi\in\mathcal{P}_{1}(\mathbb{R}^{d})\},$ (7)
where $\mathcal{E}$ denotes the negative Shannon entropy of a measure, and $\mathcal{T}(\xi,\rho)$ refers to the optimal value of equation 4. When $\rho$ meets the conditions enunciated in Cordero-Erausquin and Klartag (2015)Santambrogio showed that the minimizers of equation 7 coincide with the Gibbs distributions $\mathfrak{P}_{u}$, where $u$ are the convex functions whose moment measure is $\rho$.
3 Conjugate Moment Measures
We denote the convex function $u$ whose moment measure is $\rho$ as the moment potential of the distribution $\rho$. Additionally, we refer to the Gibbs distribution associated with such potential, $\mathfrak{P}_{u}$ as in equation 5, as the Gibbs factor of $\rho$.
3.1 Limitations of Moment Measures
The general case
By rewriting the KL divergence between $\rho$ and $\mathfrak{P}_{u^{*}}$ as follows:
$\textrm{KL}(\rho\|\mathfrak{P}_{u^{*}})=\mathcal{E}(\rho)+\int u^{*}\mathrm{d}\rho+\ln\left(\int e^{-u^{*}}\right),$
and applying this in the variational problem studied by Cordero-Erausquin and Klartag as referenced in equation 6, we obtain that the moment potentials of $\rho$ minimize the following problem:
$\min_{\begin{subarray}{c}u\in\mathcal{C}(\mathbb{R}^{d})\end{subarray}}\textrm{KL}(\rho\|\mathfrak{P}_{u^{*}})-\ln\left(\int e^{-u}\int e^{-u^{*}}\right)\,.$ (8)
The quantity $\int e^{-u}\int e^{-u^{*}}$ is intrinsic to the density $\mathfrak{P}_{u}$, and has been investigated by Ball (1986) and Artstein-Avidan et al. (2004) who demonstrated that, up to a translation, this term is maximized when $\mathfrak{P}_{u}$ is Gaussian. This reformulation reveals that the Gibbs distributions associated with the Legendre transform of $\rho$’s moment potentials, $\mathfrak{P}_{u^{*}}$, are the closest log-concave distribution to $\rho$, adjusted by the regularization term $\int e^{-u}\int e^{-u^{*}}$. As a result, the Gibbs factor of $\rho$ may differ significantly from $\rho$, especially in cases where $\rho$ is a Gaussian distribution.
Gaussian distributions. The multivariate normal distribution $\mathcal{N}(m,\Sigma)$ with $\Sigma$ non-degenerate is recovered as by $\mathfrak{P}_{v}$ with $v(x)=\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)$. When $m\neq 0$ or $\Sigma$ is degenerate, Cordero-Erausquin and Klartag show that $\mathcal{N}(m,\Sigma)$ is not a moment measure. Yet, when $\Sigma$ is non-degenerate, moment potentials of $\mathcal{N}(0,\Sigma)$ exist and have an analytical expression given by the following proposition.
Proposition 1.
Let $\rho=\mathcal{N}(0_{\mathbb{R}^{d}},\Sigma)$. If $\Sigma$ is non degenerate, the moment potential of $\rho$ is
$u_{m}(x)=\frac{1}{2}(x-m)^{T}\Sigma(x-m),$
for some $m\in\mathbb{R}^{d}$. The Gibbs factor of $\rho$ is $\mathcal{N}(m,\Sigma^{-1})$.
The proof is provided in the Appendix and relies on the fact that the OT map between two Gaussians is known in closed form Peyré et al. (2019). Intuitively this results can be interpreted as follows. When a peaked distribution $\rho$ is centered around the origin (e.g. $\Sigma\approx 0$), its corresponding moment potential is such that the image set $\nabla u$ on the support of $\mathfrak{P}_{u}$ is necessarily tightly concentrated around 00. This has the implication that $u$ is a slowly (almost constant) varying potential on the entire support of $\mathfrak{P}_{u}$. As a result, one has the (perhaps) counter-intuitive result that the more peaky $\rho$, the more spread-out $\mathfrak{P}_{u}$ must be. From this simple observation, validated experimentally in Section 3.3 we draw the intuition that a change is needed to reverse this relationship, while still retaining the interest of a measure factorization result.
Figure 1: Gibbs factor and conjugate Gibbs factor  of $\rho=\mathcal{N}\left(0,\begin{bmatrix}2&1.8\\ 1.8&2\end{bmatrix}\right)$.
3.2 Adding a Conjugate
Our main theorem is stated as follows:
Theorem 1.
Let $\rho$ be an absolutely continuous probability measure supported on a compact, convex set. Then, there exists a convex function $w$ such that
$\rho=\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}.$
Similarly to the terminology we used for moment measures, we refer to $\rho$ as the conjugate moment measure of $w$, the potential $w$ as the conjugate moment potential of $\rho$, and $\mathfrak{P}_{w}$ as $\rho$’s conjugate Gibbs factor.
Proof Sketch. The complete proof, along with the theorems used in the proof, can be found in the Appendix. We rely on the identity $\nabla w^{*}\circ\nabla w=\mathrm{id}$ that holds for strictly convex functions $w$ and show the existence of a strictly convex potential $w$ that verifies $\nabla w\,\sharp\,\rho=\mathfrak{P}_{w}$. To proceed, we first introduce some notations. Let $\mathcal{L}(\mathbb{R})$ be the set of functions mapping $\mathbb{R}^{d}$ to {$\mathbb{R}\cup\{+\infty\}$. Given a non-negligible set $\Omega\subset\mathbb{R}^{d}$ and a continuous function $v:\Omega\rightarrow\mathbb{R}$ such that $\int_{\mathbb{R}^{d}}e^{-v(z)}\mathrm{d}z$ is finite, we denote by $\mathfrak{P}_{v}^{\Omega}$ the probability measure with density
$\frac{e^{-v(x)}1_{x\in\Omega}}{\int_{\Omega}e^{-v(z)}\mathrm{d}z}\,.$
We then define, for a probability $\rho$ supported on a non-negligible set $\Omega$, the map $G_{\rho}^{\Omega}:\mathcal{L}(\mathbb{R})\rightarrow\mathcal{L}(\mathbb{R})$ which to a potential $v$ associates the Brenier potential, defined in equation 4, from $\rho$ to $\mathfrak{P}_{v}^{\Omega}$
$G_{\rho}^{\Omega}(v):=\mathfrak{B}(\rho,\mathfrak{P}_{v}^{\Omega})$
The fixed points of $G_{\rho}^{\Omega}$ are precisely the convex functions $v$ such that $\nabla v\,\sharp\,\rho=\mathfrak{P}_{v}^{\Omega}$ from which we can construct $w$ defined as $w(x)=v(x)$ for $x\in\Omega$ and $w(x)=+\infty$ elsewhere. This function $w$ is convex and verifies $\nabla w\,\sharp\,\rho=\mathfrak{P}_{w}$ (since $\rho$ is supported on $\Omega$). We rely on Schauder’s fixed point theorem to show that $G_{\rho}^{\Omega}$ admits a fixed-point. Given a Banach space $(X,\|.\|)$ and a compact, convex and nonempty set $\mathcal{M}\subset X$, Schauder’s theorem states that any continuous operator $A:\mathcal{M}\rightarrow\mathcal{M}$ has at least one fixed point. In this proof, we consider the Banach space of continuous functions over a compact set $\Omega$, denoted $\mathscr{C}(\Omega)$, equipped with the supremum norm, and establish the following theorem that directly implies Theorem 1.
Theorem 2.
Let $\rho$ be an absolutely continuous probability measure supported on a compact, convex set $\Omega\subset\mathbb{R}^{d}$. Then $G_{\rho}^{\Omega}$ admits a fixed point in
$\mathcal{M}$ =$=\{f\in\mathscr{C}(\Omega)\;\text{such that}\;\;\forall x,y\in\Omega,$
$|f(x)-f(y)|\leq R\|x-y\|_{2}\;\text{and}\;f(0_{\mathbb{R}^{d}})=0\}$
where $R$ is the radius of an euclidean ball that contains $\Omega$ i.e., $\Omega\subset\mathcal{B}(0,R)=\{x\in\mathbb{R}^{d},\|x\|_{2}\leq R\}$.
We first show that the set $\mathcal{M}$ is a non-empty, compact, convex set of $X=(\mathscr{C}(\Omega),\|\;\|_{\infty})$ by relying Arzela-Ascoli’s theorem for the compactness. We then use Brenier’s theorem that, given the absolute continuity of $\rho$, garantees the existence of the Brenier potential $\mathfrak{B}(\rho,\mathfrak{P}_{v}^{\Omega})$ and the fact that $G_{\rho}^{\Omega}$ is therefore well-defined on the set $\mathcal{M}\subset\mathscr{C}(\Omega)$. Moreover, the gradient of the obtained potential $G_{\rho}^{\Omega}(v)$ transports $\rho$ on $\mathfrak{P}_{v}^{\Omega}$ which is compactly supported on $\Omega\subset B(0,R)$. For this reason, $\nabla G_{\rho}^{\Omega}(v)$ is bounded by $R$ on $\Omega$ and the obtained potential is Lipschitz continuous with constant $R$. This permits to show that $G_{\rho}^{\Omega}(\mathcal{M})\subseteq\mathcal{M}$. To prove the continuity of $G_{\rho}^{\Omega}$ on $\mathcal{M}$, one can remark that $G_{\rho}^{\Omega}=H_{\rho}\circ F^{\Omega}$, with $F^{\Omega}(v)=\mathfrak{P}_{v}^{\Omega}$ and $H_{\rho}(\mu)=\mathfrak{B}(\rho,\mu)$, both viewed as functions from $\mathscr{C}(\Omega)$ to $\mathscr{C}(\Omega)$. To show that $F^{\Omega}$ is continuous, we use the definition of continuity in $X=(\mathscr{C}(\Omega),\|\;\|_{\infty})$ while we rely on Theorem 1.52 from Santambrogio (2015) to prove that $H_{\rho}$ is continuous. We conclude by applying Schauder’s theorem.
3.3 Illustrations
We focus on the case where $\rho$ is Gaussian to verify that our conjugate factorization has better properties. Recall that we saw in subsection 3.1 that Gibbs factors of $\mathcal{N}(0,\Sigma)$ were multivariate normal distributions with covariance matrix $\Sigma^{-1}$.
Proposition 2.
When $\Sigma$ is non degenerate, $\rho=\mathcal{N}(m,\Sigma)$ is the conjugate moment measure of
$w(x)=\frac{1}{2}(x-r)^{T}\Sigma^{-1/3}(x-r),$
whose Gibbs distribution is $\mathfrak{P}_{w}=\mathcal{N}(r,\Sigma^{1/3})$ where $r=(I_{d}+\Sigma^{1/3})^{-1}m$. Potential $w$ is the unique potential for which the Gibbs distribution is Gaussian, satisfying the condition $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho\,.$
The proof is similar to that of Proposition 1 and is given in the Appendix. While this potential $w$ may not be the only conjugate moment potential, the Gibbs distribution associated with it, $\mathcal{N}(r,\Sigma^{1/3})$, appears to be better suited to the target distribution $\rho$. This is particularly evident in the example shown in Figure 1. Additionally, it is worth noting that for the Gaussian case, our factorization still works when $\rho$ is not centered, whereas the appraoch of Cordero-Erausquin and Klartag (2015) requires $\rho$ to be mean 0.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d)
Figure 2: Comparison between the Gibbs factor $\mathfrak{P}_{u}$ and the conjugate Gibbs factor $\mathfrak{P}_{w}$ for two mixtures of 1D Gaussian distributions, $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$. The density plots overlay the (conjugate) Gibbs factor with $\bm{\rho}$ and a standard Gaussian distribution $\bm{\mathcal{N}(0,1)}$ for reference. Plots (a) and (c) show that the usual Gibbs factor’s spread is inverse to that of $\rho$. In plots (b) and (d) we see the far more suitable behavior of our conjugate moment factorization.
4 Estimating Conjugate Moment Measures from Data Samples
We first explain how to sample from $\rho$ when its conjugate moment decomposition is known using the Langevin Monte Carlo (LMC) algorithm and a conjugate solver. We then describe a method to estimate a conjugate moment potential of a distribution $\rho$ using i.i.d samples $(x_{1},\dots,x_{n})\sim\rho$. In this approach, the conjugate moment potential of $\rho$ is parameterized using an input convex neural network (ICNN) $w_{\theta}$ following the architecture proposed in (Vesseron and Cuturi, 2024). The conjugate moment potential $w_{\theta}$ is then estimated using an algorithm inspired by the fixed-point method associated to Theorem 2.
4.1 Sampling from a conjugate moment measure
In this paragraph, we suppose that we know a conjugate moment potential $w$ of $\rho$, i.e. $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. Knowing $w$, drawing samples from $\rho$ can be done by first sampling $x_{i}\sim\mathfrak{P}_{w}$ and then applying $\nabla w^{*}$ to those points as $\nabla w^{*}(x_{i})\sim\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. The LMC algorithm is a widely used method for generating samples from a smooth, log-concave density like $\mathfrak{P}_{w}$ (Roberts and Tweedie, 1996; Cheng and Bartlett, 2018; Dalalyan and Karagulyan, 2019). Starting from an initial point $x^{(0)}$, the LMC algorithm iterates according to the following update rule:
$x^{(k+1)}$ $=x^{(k)}-\gamma\nabla w(x^{(k)})+\sqrt{2\gamma}z^{(k)},\;\;z^{(k)}\sim\mathcal{N}(0,I_{d}).$
where $\gamma$ is the step size. After a warm-up period, the LMC iterates are distributed according to the log-concave distribution $\mathfrak{P}_{w}$. As for the gradient of the convex conjugate $\nabla w^{*}$, it can be efficiently estimated from $w$. By applying Danskin’s envelope theorem (1966), it follows that $\nabla w^{*}(y)$ is the solution to the following concave maximization problem,
$\nabla w^{*}(y)=\arg\sup_{x}\,\langle x,y\rangle-w(x)\,.$
This optimization problem can be solved using algorithms such as gradient ascent, (L)BFGS (Liu and Nocedal, 1989), or ADAM (Kingma and Ba, 2014). Thus, having access to a conjugate moment potential $w$ of $\rho$ enables to efficiently sample from the target distribution $\rho$. When the conjugate solver is warm-started with $x_{0}=y$, gradient steps can be interpreted as a denoising procedure of the Langevin sample.
4.2 Fixed-point Approach
For a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$, we define the map $G_{\rho}:\mathcal{L}(\mathbb{R^{d}})\rightarrow\mathcal{L}(\mathbb{R^{d}})$ which to a potential $w$ associates the Brenier potential from $\rho$ to $\mathfrak{P}_{w}$
$G_{\rho}(w):=\mathfrak{B}(\rho,\mathfrak{P}_{w})$
The fixed point of $G_{\rho}$ correspond exactly to the conjugate moment potentials of $\rho$. This observation motivates the following fixed-point iteration scheme to compute a conjugate moment potential of $\rho$:
$w_{0}:=\tfrac{1}{2}\|\cdot\|^{2};\quad\forall t\geq 1,\quad w_{t+1}:=G_{\rho}(w_{t}).$
1:  Initialize $w_{\theta}$ such that $w_{\theta}\approx\tfrac{1}{2}\|\cdot\|^{2}$
2:  while not converged do
3:     Draw $n$ i.i.d samples $x_{i}\sim\rho$
4:     Draw $y_{1},\dots,y_{n}\sim\mathfrak{P}_{w_{\theta}}$ using LMC algorithm
5:     $\mathcal{L}_{\theta}\leftarrow\frac{1}{n}\sum_{i=1}^{n}w_{\theta}(x_{i})-\frac{1}{n}\sum_{i=1}^{n}w_{\theta}(\tilde{x}(y_{i}))$
6:     Update $w_{\theta}$ with $\nabla\mathcal{L}_{\theta}$
7:  end while
Algorithm 1 Fixed point training of $w_{\theta}$
Starting from the standard Gaussian distribution $\mathfrak{P}_{w_{0}}=\mathcal{N}(0,I_{d})$, we iteratively compute the OT potential $w_{t+1}$ between the distribution $\rho$ and $\mathfrak{P}_{w_{t}}$ at iteration $t+1$. During the next iteration $t+2$, the updated distribution $\mathfrak{P}_{w_{t+1}}$ becomes the target distribution for computing the next OT map starting from the source $\rho$. We apply this methodology when the OT map is available in closed form at each step. In all other cases, we found that performing just one optimization step to estimate the OT map at each iteration was sufficient. For this reason, our final methodology relies on a single ICNN $w_{\theta}$ which serves two purposes at each iteration: generating samples from $\mathfrak{P}_{w_{\theta}}$ and being optimized by doing one step of the neural OT solver from Amos (2023) to approximate the OT map between $\rho$ and $\mathfrak{P}_{w_{\theta}}$. Algorithm 1 details the steps of the procedure, where $\tilde{x}(y_{i})$ is the estimation of $\nabla w_{\theta}^{*}(y_{i})$, computed as described in Amos (2023).
5 Experiments
We conduct preliminary experiments using the fixed-point iteration scheme described in Section 4.2, focusing on univariate $\rho$, for which Monge maps have a closed form. Following this, we estimate the conjugate moment potential associated with several 2D distributions as well as the MNIST digits distribution (LeCun et al., 2010).
5.1 Univariate distributions
In the 1D case, the OT map between density functions $\mu$ and $\nu$ is given by the formula (Peyré et al., 2019):
$\mathfrak{B}(\mu,\nu)=C_{\nu}^{-1}\circ C_{\mu}$
where $C_{\mu}:\mathbb{R}\rightarrow[0,1]$ is the cumulative distribution function associated to $\mu$, defined as: $C_{\mu}(x):=\int_{-\infty}^{x}\mathrm{d}\mu$. The quantile function $C_{\mu}^{-1}:[0,1]\rightarrow\mathbb{R}\cup\{-\infty\}$ is the pseudoinverse $C_{\mu}^{-1}(r):=\min\{x\in\mathbb{R}\cup\{-\infty\}:C_{\mu}(x)\geq r\}$. For a given distribution $\rho$ and an initial distribution $\mathfrak{P}_{w_{0}}$, the analytical form of the Brenier map enables the exact application of the fixed-point algorithm described in section 4.2. Further implementation details can be found in the Appendix. We use this methodology to compute the moment potential and the conjugate moment potential of two distinct mixtures of Gaussians. Display (a) in Figure 2 shows that the Gibbs factor of $\rho_{1}$ has a longer tail compared to the centered standard Gaussian. This aligns with theoretical results on Gaussians, which suggest that more concentrated distributions yield broader Gibbs factors. Conversely, for the second measure $\rho_{2}$ in (c), which has a larger variance, the Gibbs factor is more concentrated. By design, our conjugate Gibbs factor in (b) and (d) closely matches the shapes of $\rho_{1}$ and $\rho_{2}$, making it a more suitable prior for generative modeling applications. See also Figure 5 in the appendix, where we display $\nabla u\,\sharp\,\mathfrak{P}_{u}$ (for the moment potential $u$) and $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}$ (for the conjugate moment potential $w^{*}$) with their respective measures $\rho_{1}$ and $\rho_{2}$.
Measure $\rho$, potential $w_{\theta}$ $\nabla w_{\theta}^{*}\,\sharp\,e^{-w_{\theta}}$efer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)
Figure 3: Measure $\rho$ and level sets of its convex conjugate potential $w_{\theta}$ (top), and $\nabla w_{\theta}^{*}\,\sharp\,e^{-w_{\theta}}$ (bottom).
5.2 Generative Modeling
We focus on several 2D distributions as well as the MNIST dataset (LeCun et al., 2010). Our first step is to estimate a conjugate moment potential $w$ for these distributions using the fixed-point method outlined in Section 4.2. Following this, we generate new images based on the conjugate potential $w$ by utilizing the methodology detailed in Section 4.1. The hyperparameters used for these experiments are provided in the Appendix.
[…]
Figure 4: Learning of the conjugate moment measure. The single-step generative model $\nabla w_{\theta}$ converts highly structured data $\rho$ into a log-concave distribution $\mathfrak{P}_{w_{\theta}}$. The train distribution $\rho$ is sampled via optimization, in an iterative manner. Strict convexity of $w_{\theta}$ guarantees both (i) invertibility of $\nabla w_{\theta}$, and (ii) that Langevin dynamics recovers correctly the conjugate moment $\mathfrak{P}_{w_{\theta}}$.
2D Experiments. For the 2D distributions: (a) Circles, (b) S-curve, (c) Checkerboard, (d) Scaled-Rotated S-curve and (e) Diag-Checkerboard, the fixed-point approach proved effective. The results, shown in Figure 3, demonstrate that we successfully estimated the conjugate potential associated with these distributions, as the conjugate moment measure of the estimated potential is well aligned with the target distributions $\rho$. Additionally, in the 2D experiments (c), (d) and (e), we can clearly observe how the Gibbs factor closely follows the shape of the distribution $\rho$.
MNIST Experiment. The fixed-point method also performs well on the MNIST dataset. As shown in Figure 4, we successfully generated digits using our approach. To the best of our knowledge, this is the first instance where the MNIST distribution has been generated using an ICNN. The structure of the Gibbs distribution, compared to Gaussian noise, seems to facilitate the generative task of transporting $\mathfrak{P}_{w}$ to the MNIST distribution.
6 Conclusion
In this work, we borrow inspiration from  (Cordero-Erausquin and Klartag, 2015; Santambrogio, 2016) to define a new representation for probability measures. For any $\rho$, we establish the existence of a convex potential $w$ such that $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. We show that this representation is better suited to generative modelling, since the Gibbs factor $\mathfrak{P}_{w}$ follows more closely the original measure $\rho$, in contrast to Cordero-Erausquin and Klartag’s approach, $\nabla u\,\sharp\,\mathfrak{P}_{u}=\rho$, which results in a Gibbs factor $\mathfrak{P}_{u}$ whose spread is inversely proportional to that of $\rho$. Our conjugate measure factorization provides a single function $w$ to sample noises (using LMC, guaranteed to work with log-concave distributions) and transform these codes using $\nabla w^{*}$. We propose to parameterize the conjugate potential $w$ as an ICNN $w_{\theta}$ and rely on the OT toolbox to define a fixed-point iteration scheme to estimate $w_{\theta}$. While our method demonstrated promising results on generative modeling tasks, its performance is currently hindered by difficulties in training ICNN architecture. Future work may also explore the suitability of replacing $\mathcal{N}(0,I)$ in classic approaches with our pre-trained Gibbs factor $\mathfrak{P}_{w}$.
Acknowledgements
This work was performed using HPC resources from GENCI–IDRIS (Grant 2023-103245).
Appendix A Appendix / Supplementary Material
A.1 Proof of Proposition 1
Let us observe that $\rho$ is the moment measure of $u$ if and only if $\nabla u$ is the Monge map from $\mathfrak{P}_{u}$ to $\rho$. One can then consider quadratic functions $u_{r,\Delta}(x)=\frac{1}{2}(x-r)^{T}\Delta^{-1}(x-r)$ whose Gibbs distribution is $\mathfrak{P}_{u}=\mathcal{N}(r,\Delta)$ and use the expression (Gelbrich, 1990) of the OT map between two Gaussian distributions $\mathcal{N}(m_{A},\Sigma_{A})$ and $\mathcal{N}(m_{B},\Sigma_{B})$:
$T_{m_{A},\Sigma_{A},m_{B},\Sigma_{B}}:x\mapsto m_{B}+M(x-m_{A})$
where $M=\Sigma_{A}^{-1/2}(\Sigma_{A}^{1/2}\Sigma_{B}\Sigma_{A}^{1/2})^{1/2}\Sigma_{A}^{-1/2}$, to solve the equation
$\nabla u_{r,\Delta}(x)=T_{r,\Delta,0,\Sigma}(x)\;\;\forall x\in\mathbb{R}^{d}$
One get:
$\Delta^{-1}(x-r)=M(x-r)\;\;\text{with}\;M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$
By identification, we obtain that $\Delta^{-1}=M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$ which leads to $\Delta=\Sigma^{-1}$. It means that the functions in $\{u_{r,\Sigma^{-1}},r\in\mathbb{R}^{d}\}$ are fixed-point of $G_{\rho}$ with $\rho=\mathcal{N}(0,\Sigma)$. According to Cordero-Erausquin and Klartag (2015), the moment potential is unique up to translation, which implies that there are no other convex functions whose moment measure is $\rho$. This observation can be intuitively understood from equation 8 where the term $\ln\left(\int e^{-u}\int e^{-u^{*}}\right)$ is constant, equal to $n\ln(2\pi)$ for centered multivariate normal distributions. Among these distributions, the optimal solution involves selecting $u$ such that $\mathfrak{P}_{u^{*}}=\rho$ in order to minimize the KL divergence term.
A.2 Proof of Proposition 2
The proof is similar to the proof of Proposition 1. Given $\rho=\mathcal{N}(m,\Sigma)$, we are looking for $w$ such that $\nabla w^{*}\#\mathfrak{P}_{w}=\rho$ and $\mathfrak{P}_{w}=\mathcal{N}(r,\Delta)$. This means that $w(x)=\frac{1}{2}(x-r)^{T}\Delta^{-1}(x-r)$ and $w^{*}(x)=\frac{1}{2}x^{T}\Delta x+x^{T}r$. Moreover $\nabla w^{*}:x\rightarrow\Delta x+r$ must be the OT map between $\mathfrak{P}_{w}=\mathcal{N}(r,\Delta)$ and $\rho=\mathcal{N}(m,\Sigma)$. The OT map between these two gaussians is known in closed form and is given by $x\rightarrow m+M(x-r)$ with $M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$. Solving $\Delta=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$ gives $\Delta=\Sigma^{1/3}$ while solving $r=m-Mr$ gives $r=(I_{d}+\Sigma^{1/3})^{-1}m$. We conclude that $\rho$ is the conjugate moment measure of the unique potential $w$ among those whose Gibbs distribution is Gaussian.
A.3 Proof Theorem 2
Let $\rho$ be an absolutely continuous probability measure supported on a compact, convex set $\Omega\subset\mathbb{R}^{d}$. $\Omega$ being compact, it is bounded and there exists a real number $R$ such that $\Omega$ be included in the ball $\mathcal{B}(0,R)$. We begin by recalling Schauder’s fixed point theorem, which is central to our proof for showing that $G_{\rho}^{\Omega}$, defined as:
$G_{\rho}^{\Omega}(v):=\mathfrak{B}(\rho,\mathfrak{P}_{v}^{\Omega})$
where $\mathfrak{P}_{v}^{\Omega}$ the probability measure with density
$\frac{e^{-v(x)}1_{x\in\Omega}}{\int_{\Omega}e^{-v(z)}\mathrm{d}z}\,,$
admits a fixed-point.
Theorem (Schauder’s fixed point Theorem).
Let $(X,\|.\|)$ be a Banach space and $\mathcal{M}\subset X$ is compact, convex, and nonempty. Any continuous operator $A:\mathcal{M}\rightarrow\mathcal{M}$ has at least one fixed point.
We consider the Banach space $X=(\mathscr{C}(\Omega),\|\;\|_{\infty})$ to study the continuity of $G_{\rho}^{\Omega}$ on the set
=$\mathcal{M}=\{f\in\mathscr{C}(\Omega)\;\text{such that}\;\forall x,y\in\Omega,|f(x)-f(y)|\leq R\|x-y\|_{2}\;\text{and}\;f(0_{\mathbb{R}^{d}})=0\}.$
We prove the two following lemmas in order to use Schauder’s theorem.
Lemma 1.
The set $\mathcal{M}$ is a non-empty, compact, convex set of $X$.
Lemma 2.
$G_{\rho}^{\Omega}$ is well defined and continuous on $\mathcal{M}$ and $G_{\rho}^{\Omega}(\mathcal{M})\subseteq\mathcal{M}$.
By applying these two lemmas along with Schauder’s theorem, we conclude that $G_{\rho}^{\Omega}$ has a fixed point $v_{\text{opt}}$ in $M\subset\mathscr{C}(\Omega)$. Moreover, the absolute continuity of $\mathfrak{P}_{v_{\text{opt}}}^{\Omega}$ ensures that the optimal solution $v_{\text{opt}}$ whose gradient transports $\rho$ onto $\mathfrak{P}_{v_{\text{opt}}}^{\Omega}$ is strictly convex on the support of $\rho$, $\Omega$. Additionally, the gradient of its Legendre transform $\nabla v^{*}_{\text{opt}}$ is the OT map from $\mathfrak{P}_{v_{\text{opt}}}^{\Omega}$ to $\rho$.
A.3.1 Proof of Lemma 1.
$\mathcal{M}$ is non-empty
00 is included in $\mathcal{M}$ which is therefore not empty.
$\mathcal{M}$ is convex
Let $f$ and $g$ functions of $\mathcal{M}$ and $\lambda\in[0,1]$. We have that $\lambda f+(1-\lambda)g\in\mathscr{C}(\Omega)$ and that $\lambda f(0)+(1-\lambda)g(0)=0$. Let $x\in\Omega$ and $y\in\Omega$, then :
$|(\lambda f(x)+(1-\lambda)g(x))-(\lambda f(y)+(1-\lambda)g(y))|$ $\leq\lambda|f(x)-f(y)|+(1-\lambda)|g(x)-g(y)|$
$\leq\lambda R\|x-y\|_{2}+(1-\lambda)R\|x-y\|_{2}=R\|x-y\|_{2}$
where, for the first inequality, we rely on the fact that the absolute value verifies the triangular inequality and for the second inequality, we use that $f$ and $g$ belong to $\mathcal{M}$ and are therefore $R$-lipschtiz.
$\mathcal{M}$ is compact
We rely on Arzela-Ascoli theorem which is recalled below to show that the set $\mathcal{M}$ is compact. We also recall the definitions of equicontinuity and equiboundness.
Definition.
A family of functions $\mathcal{F}\subset\mathscr{C}(\Omega)$ is equibounded if there exists a constant $C_{0}$ such that for all $x\in\Omega$ and $f\in\mathcal{F}$, we have $|f(x)|\leq C_{0}$.
Definition.
A family of functions $\mathcal{F}\subset\mathscr{C}(\Omega)$ is equicontinous if for all $\varepsilon>0$, there exists $\delta$ such that for $x,y\in\Omega$,
$\|x-y\|<\delta\implies\forall f\in\mathcal{F}\;|f(x)-f(y)|<\varepsilon$
Theorem (Arzela-Ascoli).
Let $(K,d)$ a compact space, a family of functions $\mathcal{F}\subset\mathscr{C}(K)$ is relatively compact if and only if $\mathcal{F}$ is equibounded and equicontinuous.
The family of functions $\mathcal{M}$ is equibounded. In fact, for $f\in M$ and $x\in\Omega$, we have that:
$|f(x)|=|f(x)-f(0)|\leq R\|x-0\|_{2}\leq R^{2}$
The first equality comes from the fact that $f\in M$, therefore $f(0)=0$, the first inequality uses the fact that $f$ is Lipschitz continuous with constant $R$ while the last one is due to the fact that $x\in\Omega\subset B(0,R)$.
The family of functions $\mathcal{M}$ is equicontinuous. For $\varepsilon>0$, we define $\delta=\frac{\varepsilon}{R}$. Let $f\in\mathcal{M}$ and $x,y\in\Omega$ such that $\|x-y\|_{2}<\delta$ then because $f$ is Lipschitz on $\Omega$, we have that
$|f(x)-f(y)|\leq R\|x-y\|_{2}<R\delta=\varepsilon\,.$
Then, according to Arzela-Ascoli theorem, $\mathcal{M}$ is relatively compact i.e. its closure is compact.
$\mathcal{M}$ is closed in $X=(\mathscr{C}(\Omega),\|\;\|_{\infty})$
Let $(f_{n})$ be a sequence of functions in $\mathcal{M}$ that converges uniformly to a limiting function $f$. We have that $f\in\mathscr{C}(\Omega)$ which is closed. Moreover, because $(f_{n})$ converges uniformly to $f$, it also converges pointwise and in particular $f_{n}(0)=0\rightarrow 0=f(0)$. Finally, because $\forall n$, $f_{n}$ is Lipschitz continuous, we have
$|f_{n}(x)-f_{n}(y)|\leq R\|x-y\|_{2}\;\forall n$
By taking the limit when $n\rightarrow+\infty$ we get that:
$|f(x)-f(y)|\leq R\|x-y\|_{2}$
and $f$ is Lipschitz continuous with constant $R$ and $f\in\mathcal{M}$. This proves that $\mathcal{M}$ is closed and that $\mathcal{M}$ is therefore compact.
A.3.2 Proof of Lemma 2.
$G_{\rho}^{\Omega}$ is well-defined on $\mathcal{M}$
For a given $v\in\mathcal{M}$, the fact that $v\in\mathscr{C}(\Omega)$ and that $\Omega$ is non-negligible (because $\rho$ is an absolutely continuous probability measure supported on $\Omega$) ensure the measure $\mathfrak{P}_{v}^{\Omega}$ to be well-defined. Using Brenier’s theorem, the absolute continuity of $\rho$ garantees the existence of the Brenier potential $\mathfrak{B}(\rho,\mathfrak{P}_{v}^{\Omega})$ and the fact that $G_{\rho}^{\Omega}$ is therefore well-defined on the set $\mathcal{M}\subset\mathscr{C}(\Omega)$.
$G_{\rho}^{\Omega}(\mathcal{M})\subset\mathcal{M}$
For a given $v\in\mathcal{M}$, the Brenier potential $G_{\rho}^{\Omega}(v)$ verifies $G_{\rho}^{\Omega}(v)(0)=0$ by definition. As a convex function which takes real values on $\Omega$, $G_{\rho}^{\Omega}(v)$ is continuous on $\Omega$ and $G_{\rho}^{\Omega}(v)\in\mathscr{C}(\Omega)$. Moreover, the gradient of the obtained potential $\nabla G_{\rho}^{\Omega}(v)$ transports $\rho$ on $\mathfrak{P}_{v}^{\Omega}$. The probability measures $\mathfrak{P}_{v}^{\Omega}$ and $\rho$ are both compactly supported on $\Omega\subset B(0,R)$ which implies that the support of the optimal transport plan $\gamma=(i_{d},\nabla G_{\rho}^{\Omega}(v))_{\#}\rho$ is included in $\Omega\times\Omega$. $\Omega$ being the support of $\rho$, we have that for each $x\in\Omega$, $\nabla G_{\rho}^{\Omega}(v)(x)\in\Omega\subset B(0,R)$ and in particular $\|\nabla G_{\rho}^{\Omega}(v)(x)\|_{2}\leq R$. $\Omega$ being convex and $G_{\rho}^{\Omega}(v)$ being differentiable in the interior of $\Omega$, we can apply the mean value theorem to deduce that:
$\forall x,y\in\Omega\;|G_{\rho}^{\Omega}(v)(x)-G_{\rho}^{\Omega}(v)(y)|\leq\sup_{z\in\Omega}\|\nabla G_{\rho}^{\Omega}(v)(z)\|_{2}\|x-y\|_{2}\leq R\|x-y\|_{2}$
$G_{\rho}^{\Omega}(v)$ is therefore Lipschitz continuous with constant $R$ and it belongs to the set $\mathcal{M}$.
$G_{\rho}^{\Omega}$ is continuous on $\mathcal{M}$
To prove the continuity of $G_{\rho}^{\Omega}$ on $\mathcal{M}$, one can remark that $G_{\rho}^{\Omega}=H_{\rho}\circ F^{\Omega}$, with
$F^{\Omega}:\;\mathcal{M}$ $\longrightarrow\mathcal{N}$
$v$ $\longrightarrow\mathfrak{P}_{v}^{\Omega}$
and
$H_{\rho}:\;\mathcal{N}$ $\longrightarrow\mathscr{C}(\Omega)$
$\mu$ $\longrightarrow\mathfrak{B}(\rho,\mu)$
with $\mathcal{N}=\{\mathfrak{P}_{v}^{\Omega},v\in\mathscr{C}(\Omega)\}\subset\mathscr{C}(\Omega)$. We show that both function $F^{\Omega}$ and $H_{\rho}$ are continuous in $X=(\mathscr{C}(\Omega),\|\;\|_{\infty})$. Let $v\in\mathscr{C}(\Omega)$ and $(v_{n})$ a familly of functions of $\mathscr{C}(\Omega)$ that converges uniformly to $v$. We will show that the family $\left(F^{\Omega}(v_{n})\right)$ converges uniformly to $F^{\Omega}(v)$. Let $x$ in $\Omega$ and $n\in\mathbb{N}$, we have:
$|F^{\Omega}(v_{n})(x)-F^{\Omega}(v)(x)|$ $=\Bigg{|}\frac{e^{-v_{n}(x)}}{\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z}-\frac{e^{-v(x)}}{\int_{\Omega}e^{-v(z)}\mathrm{d}z}\Bigg{|}$
$=\Bigg{|}\frac{\left(e^{-v_{n}(x)}\int_{\Omega}e^{-v(z)}\mathrm{d}z\right)-\left(e^{-v(x)}\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\right)}{\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\int_{\Omega}e^{-v(z)}\mathrm{d}z}\Bigg{|}$
We will use the following lemma:
Lemma 3.
If $(v_{n})$ converges uniformly to $v$ in $\mathscr{C}(\Omega)$ with $\Omega$ compact then $e^{-v_{n}}$ converges uniformly to $e^{-v}$ in $\mathscr{C}(\Omega)$.
Proof of Lemma 3..
Let us denote by $f$ the real function $f:x\longrightarrow e^{-x}$. $f$ is continuous on $\mathbb{R}$. Because $v\in\mathscr{C}(\Omega)$ and $\Omega$ compact, we have, by Heine’s theorem, that $v(\Omega)\subset\mathbb{R}$ is compact and therefore bounded. There exists $a$ and $b$ such that $v(\Omega)\subseteq[a,b]$. Moreover because $(v_{n})$ converges uniformly to $v$, there exists $N$ such that $n\geq N\implies v_{n}(\Omega)\subseteq[a-1,b+1]$. $f$ being continuous on $\mathbb{R}$, it is uniformly continuous on the interval $[a-1,b+1]$ according to Heine’s theorem. We deduce that $(f(v_{n}))$ converges uniformly to $f(v)$.∎
On the other hand, the function $g:t\longrightarrow e^{-v(t)}$ is continuous on the compact $\Omega$ as a composition of continuous functions. By Heine’s theorem, the image of $\Omega$ by $g$ is a compact and there exists $c,\ell\in\mathbb{R}^{+*}$ such that $g(\Omega)\subset[c,\ell]$ with $c>0$. Because $e^{-v_{n}}$ converges uniformly to $e^{-v}$ in $\mathscr{C}(\Omega)$, there exists $N_{0}$ such that $n\geq N_{0}\implies e^{-v_{n}}(z)\geq\frac{c}{2}\;\forall z\in\Omega$. As a consequence, we have for $n\geq N_{0}$:
$\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\int_{\Omega}e^{-v(z)}\mathrm{d}z$ $\geq\int_{\Omega}\frac{c}{2}\mathrm{d}z\int_{\Omega}c\mathrm{d}z=\frac{c^{2}}{2}|\Omega|^{2}$
with $|\Omega|$ the lebesgue measure of the compact $\Omega$. $\Omega$ being non-negligible, we have $|\Omega|>0$. Let $C_{0}=\frac{2}{c^{2}|\Omega|^{2}}$, we have shown that for $n\geq N_{0}$ we had:
$\Bigg{|}\frac{1}{\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\int_{\Omega}e^{-v(z)}\mathrm{d}z}\Bigg{|}<C_{0}$
Furthermore, one notes that:
$\Bigg{|}\left(e^{-v_{n}(x)}\int_{\Omega}e^{-v(z)}\mathrm{d}z\right)-\left(e^{-v(x)}\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\right)\Bigg{|}$
$=\Bigg{|}\left(e^{-v_{n}(x)}\int_{\Omega}e^{-v(z)}\mathrm{d}z\right)-\left(e^{-v(x)}\int_{\Omega}e^{-v(z)}\mathrm{d}z\right)+\left(e^{-v(x)}\int_{\Omega}e^{-v(z)}\mathrm{d}z\right)-\left(e^{-v(x)}\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\right)\Bigg{|}$
$\leq\Bigg{|}\left(e^{-v_{n}(x)}-e^{-v(x)}\right)\int_{\Omega}e^{-v(z)}\mathrm{d}z\Bigg{|}+\Bigg{|}e^{-v(x)}\left(\int_{\Omega}e^{-v(z)}\mathrm{d}z-\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\right)\Bigg{|}$
$\leq\Big{|}e^{-v_{n}(x)}-e^{-v(x)}\Big{|}\ell|\Omega|+\ell\Bigg{|}\int_{\Omega}e^{-v(z)}\mathrm{d}z-\int_{\Omega}e^{-v_{n}(z)}\mathrm{d}z\Bigg{|}$
$\leq\big{\|}e^{-v_{n}}-e^{-v}\big{\|}_{\infty}\ell|\Omega|+\ell\int_{\Omega}\Big{|}e^{-v(z)}-e^{-v_{n}(z)}\Big{|}\mathrm{d}z$
$\leq 2\ell|\Omega|\big{\|}e^{-v_{n}}-e^{-v}\big{\|}_{\infty}$
Until now, we have shown that for $n\geq N_{0}$ we had, for $x\in\Omega$:
$|F^{\Omega}(v_{n})(x)-F^{\Omega}(v)(x)|\leq C_{0}2\ell|\Omega|\big{\|}e^{-v_{n}}-e^{-v}\big{\|}_{\infty}$
Let $\varepsilon>0$, let us define $\varepsilon_{0}=\frac{\varepsilon}{C_{0}2\ell|\Omega|}>0$. Because $(e^{-v_{n}})$ converges uniformly to $e^{-v}$, there exists $N_{1}$ such that $n\geq N_{1}\implies\|e^{-v_{n}}-e^{-v}\big{\|}_{\infty}<\varepsilon_{0}$. Let us denote $N=\max(N_{0},N_{1})$, we have that for $n\geq N$,
$|F^{\Omega}(v_{n})(x)-F^{\Omega}(v)(x)|\leq\varepsilon$
Because this is true for any $x\in\Omega$, we have $\big{\|}F^{\Omega}(v_{n})-F^{\Omega}(v)\big{\|}_{\infty}\leq\varepsilon$. We just proved that $F^{\Omega}$ is continuous at $v$ for any $v\in\mathscr{C}(\Omega)$ ie $F^{\Omega}$ is continuous on $\mathscr{C}(\Omega)$ and in particular it is continuous on $\mathcal{M}\subset\mathscr{C}(\Omega)$.
We will now show that $H_{\rho}$ is continuous on $\mathcal{N}$. Let $\mu\in\mathcal{N}$ and let $(\mu_{n})$ a family of densities from $\mathcal{N}$ that converges uniformly to $\mu$. For $n\in\mathbb{N}$, we denote by $\mathcal{P}(\mu_{n})$ and $\mathcal{P}(\mu)$ the probability measures associated to the respective densities $\mu_{n}$ and $\mu$. Because $(\mu_{n})$ converges uniformly to $\mu$, we have that $\mathcal{P}(\mu_{n})$ converges in distribution to $\mathcal{P}(\mu)$. Then according to Santambrogio (2015) (Theorem 1.52), the family of Kantorovich potentials from $\rho$ to $\mathcal{P}(\mu_{n})$ with value 00 in 00 that we denote $(f^{\star}_{n})$ converges uniformly to the Kantorovich potential $f^{\star}$ from $\rho$ to $\mathcal{P}(\mu)$ whose value is 00 in 00. Therefore, the family of Brenier potentials $(H_{\rho}(\mu_{n}):t\rightarrow\tfrac{1}{2}\|t\|^{2}-f^{\star}_{n}(t))$ converges uniformly to $H_{\rho}(\mu)=\tfrac{1}{2}\|\cdot\|^{2}-f^{\star}$. We conclude that $H_{\rho}$ is continuous on $\mathcal{N}$ and by composition of continuous functions $G_{\rho}^{\Omega}$ is continuous on $\mathcal{M}$.
A.4 1D Experiments
We conducted two experiments, with $\rho$ defined as a mixture of Gaussian distributions. In the first experiment, $\rho$ is a mixture of four Gaussians: $\mathcal{N}(-0.1,0.07)$, $\mathcal{N}(0.3,0.1)$, $\mathcal{N}(0.3,0.1)$, $\mathcal{N}(0.7,0.15)$ with mixture weights $\frac{1}{7},\frac{3}{7},\frac{2}{7},\frac{1}{7}$, respectively. In the second experiment, $\rho$ is a mixture of the three Gaussians $\mathcal{N}(-0.8,0.4)$, $\mathcal{N}(1.5,0.6)$, $\mathcal{N}(3,0.5)$, with weights $\frac{1}{2},\frac{1}{3},\frac{1}{6}$. To approximate these mixtures, we used 400,000 samples and estimated their histograms using 100,000 bins. This allowed us to compute the cumulative distribution and quantile functions using numpy’s cumsum and quantile functions, which were essential for estimating the OT map at each iteration of the fixed-point algorithm. The algorithm converged after 300 iterations. Figure 5 illustrates that, in both experiments, the moment measures $\nabla u\,\sharp\,\mathfrak{P}_{u}$ and the conjugate moment measures $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}$ of the computed potentials perfectly align with $\rho_{i}$ confirming that the fixed-point algorithm successfully identified moment potentials $u$ and conjugate moment potentials $w$ of $\rho_{i}$.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Convergence of the fixed-point algorithm.
A.5 Hyperparameters and infrastructure
For the MNIST experiment, we used MNIST dataset (LeCun et al., 2010) under the terms of the Creative Commons Attribution-Share Alike 3.0 license. In this experiment, we convolve the MNIST images with Gaussian noise of standard deviation $0.3$ to ensure that the OT map from the MNIST distribution to the log-concave distribution $\mathfrak{P}_{w}$ exists, as this is guaranteed when the starting distribution is absolutely continuous (Santambrogio, 2015). This effect is visible in Figure 4.
All runs have been made on single GPU V100 or A100-40GB. All the experiments in Figure 3 were generated using a common set of hyperparameters, detailed in Table 1. The number of particles used (i.e., batch size) and the chosen step size $\gamma$ in the LMC process vary across experiments. For the 2D distributions Circles, S-curve, and Scaled-Rotated S-curve, we used 1024 particles, while for Checkerboard and Diag-Checkerboard, a larger number of particles (4096) was used to ensure the stability of the algorithm. The step size $\gamma$ used in the LMC process depends on the shape of the log concave distribution $\mathfrak{P}_{w_{\theta}^{*}}$. Specifically, we used a step size of 5e-5 for Circles, 1e-4 for S-curve and Scaled-Rotated S-curve, and 1e-3 for Checkerboard and Diag-Checkerboard.
The hyperparameters used for the MNIST experiments are given in Table 2. To determine these hyperparameters, we used commonly applied architectures for ICNNs and conducted a grid search over $\gamma$ testing values 1e-3, 5e-4, 1e-4, 5e-5 and 1e-5, as well as for the learning rate of $w_{\theta}$, considering values 1e-3, 5e-4 and 1e-4. We adopt the approach of Amos (2023), where the cost of computing the conjugate at each training step is amortized by initializing the conjugate solver with the predictions of an MLP. This MLP is trained via regression on the outputs of the conjugate solver. The MLP consists of three hidden layers with 128 units each for the 2D experiments and three hidden layers with 256 units each for the MNIST experiment. It is trained using the Adam optimizer with default parameters, and a fixed learning rate of 1e-4 for the 2D experiments and 5e-4 for the MNIST experiment. It is important to note that the number of LMC steps listed in the tables refers to the steps taken to sample from $\mathfrak{P}_{w_{\theta}}$ starting from uniform noise, as used to generate the figures. During training, however, we reuse the particles sampled from the previous gradient step and apply $200$ LMC steps to these particles to form the new batch.
Hyperparameter Value
$w_{\theta}$ architecture dense
$2\shortrightarrow 128\shortrightarrow 128\shortrightarrow 128\shortrightarrow 128\shortrightarrow 128\shortrightarrow 1$
ELU activation functions
$w_{\theta}$ optimizer Adam
step size $=0.0001$
$\beta_{1}=0.5$
$\beta_{2}=0.5$
Conjugate solver $w^{*}_{\theta}$ Adam
step size with cosine decay schedule
init value = 0.1
alpha = 1e-4
$\beta_{1}=0.9$
$\beta_{2}=0.999$
100 iterations
Langevin sampling number of steps $=6000$ (=$200$ during training)
number of gradient steps 700,000
Table 1: Hyperparameters for the 2D experiments of Figure 3.
Hyperparameter Value
$w_{\theta}$ architecture dense
$2\shortrightarrow 512\shortrightarrow 512\shortrightarrow 512\shortrightarrow 512\shortrightarrow 512\shortrightarrow 1$
ELU activation functions
$w_{\theta}$ optimizer Adam
step size $=0.0005$
$\beta_{1}=0.5$
$\beta_{2}=0.5$
Conjugate solver $w^{*}_{\theta}$ Adam
step size with cosine decay schedule
init value = 0.1
alpha = 1e-4
$\beta_{1}=0.9$
$\beta_{2}=0.999$
100 iterations
Langevin sampling step size $\gamma=0.001$
number of steps $=6000$ (=$200$ during training)
number of gradient steps 100,000
batch size 512
Table 2: Hyperparameters for the MNIST experiment of Figure 4.